# Coherent errors in stabilizer codes caused by quasistatic phase damping

### v1

@@ VI Results for the distance-3 surface code, and their relation to experiments
We now connect these results to experimentally relevant error parameters, focusing on semiconductor spin qubits, to estimate the expected performance of a small surface code experiment based on currently available spin qubit hardware. We remark that the first error correction experiments with semiconductor spin qubits realized the phase-flip repetition code [30, 31] with a single round of measurement and without feedforward; however, the surface code has yet to be realized. The solid black curve in Fig. 6 shows our estimate for the physical and readout error rates, based on recent experimental data [32] obtained for spin qubit readout in a double quantum dot. Points of the curve correspond to different measurement integration times, as we explain below. The black curve lies inside the break-even region (green), indicating that scaling up the hardware with the same high-quality components could make useful surface-code-based quantum error correction attainable with semiconductor spin qubits.
@@
The solid black curve in Fig. 6 shows the parametric dependence of $q$ and $p$ on $T_{\text{meas}}$, based on the considerations and numerical factors of the previous two paragraphs, and inhomogeneous dephasing time value [32] of $T_{2}^{\ast}=10\,\mathrm{\mu s}$. The black curve lies in the green region, where the logical error rate is lower than the physical error rate. Based on these considerations, we conclude that although it is challenging to perform a 9-qubit surface code experiment, semiconductor spin qubits have the potential to achieve break-even in the near future. In experiments, where the readout is significantly slower than in the one we cited, it could be fruitful to convert high readout fidelity to shorter readout time duration, since the surface code is more robust against readout errors than coherent errors. Note that the above considerations can be significantly improved with the goal of a more realistic description of quantum memory experiments, e.g., by taking into account gate errors both on data and measure qubits, etc.
Finally, we also relate our simulation results to superconducting quantum computing platforms. On the one hand, transmons are designed to be protected from dephasing [33], and hence their decoherence is often dominated by relaxation, which is excluded from our model. On the other hand, alternative superconducting qubits have been developed [34, 35, 36, 37], in which qubit relaxation is suppressed and pure dephasing due to charge noise or flux noise is the dominant decoherence mechanism. (For an informative summary, see Ref. [38], and Table 1 therein.) For such novel superconducting qubits, the quasistatic phase damping model is a good starting point. Therefore, it is an interesting and relevant future task to assess the performance of quantum error correction based on such qubits, utilizing the efficient simulation methodology (fermionic linear optics) applied in this paper.
@@ VII Conclusions
We have evaluated the effects of our error model on the distance-3 surface code, and established the break-even boundary line for this code on the parameter plane of the physical error rate and the readout error rate. Finally, we have discussed the relevance of our results to solid-state qubit platforms, including semiconductor spin qubits and novel superconducting qubits. Our results are expected to provide useful design guidelines for future experiments based on these quantum computing platforms.
@@ Appendix A Calculation details for the 2-qubit repetition code for two cycles of error detection
# $\boldsymbol{E}$ $\hat{E}(\boldsymbol{E})$ $\boldsymbol{s}$ $P(\boldsymbol{E})$ $\tilde{P}(\boldsymbol{E})$

### v3

@@ VI Results for the distance-3 surface code, and their relation to experiments
For physical error rates much below the threshold, $p\ll p_{\text{th}}\approx 2.85\%$, numerical simulations could only provide accurate results by significantly increasing the number of samples which is not feasible. However, it would also be interesting to see how the surface code performs in this regime. For the distance-3 surface code, we could carry out analytic calculations using a MWPM-based look-up table decoder, setting the weights of ’spacelike’ and ’timelike’ edges equal. For $3$ rounds of error correction with readout error rate equal to physical error rate, $q=p$, we find that the logical error rate up to leading order is given by
$p_{L}^{(d=3)}\approx 118p^{2}.$ (57)
Interestingly, this expression is the same for independent phase-flip and readout errors. Thus, for small error rates, we expect similar behavior for the two models.
@@
We now connect these results to experimentally relevant error parameters, focusing on semiconductor spin qubits, to estimate the expected performance of a small surface code experiment based on currently available spin qubit hardware. Our considerations expand upon earlier theoretical quantum error correction studies focusing on spin qubits [42, 43, 44, 45, 46, 47, 48]. We remark that the first error correction experiments with semiconductor spin qubits realized the phase-flip repetition code [49, 50] with a single round of measurement and without feedforward; however, the surface code has yet to be realized.
The solid black curve in Fig. 6 shows our estimate for the physical and readout error rates, based on recent experimental data [51] obtained for spin qubit readout in a double quantum dot. Points of the curve correspond to different measurement integration times, as we explain below. The black curve lies inside the break-even region (green), indicating that scaling up the hardware with the same high-quality components could make useful surface-code-based quantum error correction attainable with semiconductor spin qubits.
@@
The solid black curve in Fig. 6 shows the parametric dependence of $q$ and $p$ on $T_{\text{meas}}$, based on the considerations and numerical factors of the previous two paragraphs, and inhomogeneous dephasing time value [51] of $T_{2}^{\ast}=10\,\mathrm{\mu s}$. The black curve lies in the green region, where the logical error rate is lower than the physical error rate. Based on these considerations, we conclude that although it is challenging to perform a 9-qubit surface code experiment, semiconductor spin qubits have the potential to achieve break-even in the near future. In experiments, where the readout is significantly slower than in the one we cited, it could be fruitful to convert high readout fidelity to shorter readout time duration [46], since the surface code is more robust against readout errors than coherent errors. Note that the above considerations can be significantly improved with the goal of a more realistic description of quantum memory experiments, e.g., by taking into account gate errors both on data and measure qubits, etc.
Finally, we also relate our simulation results to superconducting quantum computing platforms. On the one hand, transmons are designed to be protected from dephasing [52], and hence their decoherence is often dominated by relaxation, which is excluded from our model. On the other hand, alternative superconducting qubits have been developed [53, 54, 55, 56], in which qubit relaxation is suppressed and pure dephasing due to charge noise or flux noise is the dominant decoherence mechanism. (For an informative summary, see Ref. [57], and Table 1 therein.) For such superconducting qubits, the quasistatic phase damping model is a good starting point. Therefore, it is an interesting and relevant future task to assess the performance of quantum error correction based on such qubits, utilizing the efficient simulation methodology (fermionic linear optics) applied in this paper.
@@ VII Conclusions
We have evaluated the effects of our error model on the distance-3 surface code, and established the break-even boundary line for this code on the parameter plane of the physical error rate and the readout error rate. Finally, we have discussed the relevance of our results to solid-state qubit platforms, including semiconductor spin qubits and superconducting qubits. Our results are expected to provide useful design guidelines for future experiments based on these quantum computing platforms.
We note that a possible extension of our error model is the combination of quasistatic phase damping and random Pauli errors. Combining random phase-flip errors with quasistatic phase damping, as considered in Ref. [58], can be an interesting future direction. Also including random Pauli $X$ and $Y$ errors, as, e.g., in depolarizing noise, however, is unfortunately beyond the scope of our approach, due to the constraints of the FLO simulation.
@@ Appendix A Calculation details for the 2-qubit repetition code for two cycles of error detection
Here, we list the 16 cycle-resolved error scenarios and the corresponding merged error operators for the 2-qubit repetition code, for two cycles of error detection. For comparison, we provide the probabilities obtained from a phase-flip error model alongside the coherent error amplitudes for each error scenario, see Table 1.
No. $\boldsymbol{E}$ $\hat{E}(\boldsymbol{E})$ $\boldsymbol{s}$ $P(\boldsymbol{E})$ $\tilde{P}(\boldsymbol{E})$